# AlGaN photonic device platform

### v1

@@ V Tunable Microring Resonators
We fabricate microring resonators, incorporating radii of 50 m and top widths of 580 nm, and coupled them to a racetrack geometry waveguide, cleaved at both ends, as depicted in Fig. 4(a). The sample is mounted on a 3D piezoelectric nanopositioner (Physik Instrumente, NanoCube) and characterized with a custom-made inverted microscope, where the same objective (Zeiss, LD EC Epiplan-Neofluar, 20X, NA = 0,22) is used to launch the laser light into one waveguide facet and collect it from the other facet. The alignment of the incoupling light is monitored with an EMCCD camera (Andor, iXon Ultra 897) and optimized by maximizing the intensity of the outcoupled signal by adjusting both sample position and laser incoupling angle. It should be noted that the diffraction-limited spot of the laser focused on the waveguide facet is significantly larger than the size of the mode supported by the waveguide. Therefore, we cannot determine the effective incoupled power. Propagation losses are then inferred by measuring the intrinsic quality factor of the microring resonators. To reach critical coupling, which allows for the estimation of resonator losses, we fabricated several microring resonators coupled to racetrack waveguides with varied coupling gaps, ranging from 100-325 nm, as displayed in Fig. 4(c). We measure the transmission of a tunable 785 nm laser (Toptica, DFB PRO centered at 784.6 nm, MHF over 2 nm) and use a linear film polarizer (Thorlabs, LPVIS) to excite and collect TE or TM modes selectively. Fig. 4(b) illustrates the typical transmission as a function of the laser’s frequency, with a baseline primarily attributed to the laser power modulation, which is typical for DFB. This modulation can be described with a quadratic baseline from which the resonant peaks can be clearly discriminated and fitted with a Lorentzian function, as indicated in red in Fig. 4(b). We extrapolated the resonance widths at the critical coupling, i.e., zero transmittance. We summarize key parameters such as the critical coupling FWHM, free spectral ranges (FSRs), and intrinsic Q-factors for TE and TM modes in Table 1.
We mount the chips on a custom-made temperature-controlled stage to test the temperature tunability of the TE and TM modes resonance frequencies in the range of 20-30 C. Fig. 4(c) depicts the shift of a resonant peak of the TE mode (coupling gap=100nm) for different temperatures. In the investigated temperature range, the central frequency of the resonance changes linearly with temperature. The corresponding temperature coefficients are 5.7 and 6.2 GHz/K for TE and TM modes. These coefficients correspond to a thermo-optic coefficient for AlGaN, which is comparable to the previously reported values for AlN (2.94 10 K) and GaN (7.01 10 K) [49] at 785 nm and 300K. We also determine the anisotropic refractive indices of the AlN and AlGaN layers by analyzing the FSRs of rings with different widths and etch depths, as explained in the supplementary material. The thermo-optic coefficient and effective index results for TM and TE modes are summarized in Table 1.
@@
| at 632 nm | 2.14 | 2.20 |
| at 632 nm | 2.07 | 2.11 |
VI Directional Couplers
@@
Directional couplers are essential components in photonic integrated circuits. They enable the transfer of electromagnetic energy between two or more waveguides that are in proximity. They can be combined to form optical switches or Mach-Zehnder interferometers, which are the fundamental constituent of the majority of on-chip quantum technology applications [50, 51]. In our study, we develop AlGaN directional couplers linked to racetrack waveguides, as depicted in Fig. 5(a). We vary coupling gaps and lengths, as illustrated in Fig. 5(b). By injecting a 632 nm laser into one of the input ports, we measure the output intensity ratio: . Our observations revealed a direct linear correlation between the coupling length and , with linear fitting of the data providing the coupling constants. For coupling gaps of 230 nm and 330 nm, the coupling constants were determined to be 0.15 cm and 0.066 cm , respectively. The observed uncertainties are primarily attributed to reflections at the input and output facets, which can be minimized by applying an anti-reflection coating. The well-functioning directional couplers show that our AlGaN platform is well-suitable for building the required components of complex integrated photonic circuits. Further improvement is to be expected by optimizing the etching processes towards vertical sidewalls.
VII Discussion
@@
Our investigation of AlGaN as a material platform for photonic integrated circuits allowed us to demonstrate the required core functional devices. By epitaxially growing AlGaN on HTA-AlN templates on sapphire substrates, we alleviate the problems related to crystal defects and dislocations at the sapphire-AlN interface. Our measured losses align with those documented in the literature concerning aluminum nitride on sapphire waveguides [52, 24, 53], even though we did not employ a cladding material, which is expected to diminish scattering loss even more. Additionally, we observed a polarization rotation effect within the rings, which might effectively increase the measured loss; however, this could be mitigated by optimizing waveguide geometry. Moreover, by measuring the thermo-optic coefficient and anisotropic refractive indices of the waveguides, we have gained insight into parameters relevant to applications such as nonlinear photon-pair generation and on-chip nonlinear conversion. The birefringent refractive index opens avenues for exploration, especially in its utilization for phase matching of TE and TM modes. Furthermore, the surface roughness of the epitaxial material is sufficiently low to integrate superconducting single-photon detectors. In summary, our work represents a significant step forward in using AlGaN for integrated photonics, offering a scalable and high-quality alternative to established materials.
@@ Acknowledgements.
We thank Karina Ickert, Nico Sluka, Joost Wartena, Adrian Runge, Mohammad Mohammadi, and Dominik Sudau for their assistance in thin-film depositions; Kevin Kunkel, Natalia Sabelfeld, Kai Gehrke, and Andreas Renkewitz for performing plasma etching processes; Alexander Külberg and Felix Eiche for laser cutting and marking; Uwe Spengler and Christine Münnich for cleaving, Olaf Krüger, Ralph-Stephan Unger and Ina Ostermay for fruitful discussions.References
- Shastri et al. [2021] B. J. Shastri, A. N. Tait, T. Ferreira de Lima, W. H. P. Pernice, H. Bhaskaran, C. D. Wright, and P. R. Prucnal, Photonics for artificial intelligence and neuromorphic computing, Nature Photonics 15, 102–114 (2021).
- Xavier et al. [2021] J. Xavier, D. Yu, C. Jones, E. Zossimova, and F. Vollmer, Quantum nanophotonic and nanoplasmonic sensing: towards quantum optical bioscience laboratories on chip, Nanophotonics 10, 1387 (2021).
- Knaut et al. [2023] C. M. Knaut, A. Suleymanzade, Y.-C. Wei, D. R. Assumpcao, P.-J. Stas, Y. Q. Huan, B. Machielse, E. N. Knall, M. Sutula, G. Baranes, N. Sinclair, C. De-Eknamkul, D. S. Levonian, M. K. Bhaskar, H. Park, M. Lončar, and M. D. Lukin, Entanglement of Nanophotonic Quantum Memory Nodes in a Telecommunication Network, (2023).
- Zhang et al. [2019] G. Zhang, J. Y. Haw, H. Cai, F. Xu, S. M. Assad, J. F. Fitzsimons, X. Zhou, Y. Zhang, S. Yu, J. Wu, W. Ser, L. C. Kwek, and A. Q. Liu, An integrated silicon photonic chip platform for continuous-variable quantum key distribution, Nature Photonics 13, 839–842 (2019).
- Liu et al. [2019] Y. Liu, K. Xu, S. Wang, W. Shen, H. Xie, Y. Wang, S. Xiao, Y. Yao, J. Du, Z. He, and Q. Song, Arbitrarily routed mode-division multiplexed photonic circuits for dense integration, Nature Communications 10, 10.1038/s41467-019-11196-8 (2019).
- Zeng et al. [2021] Z. Zeng, L. Lu, P. He, D. Liu, and M. Zhang, Demonstration of dual-mode photonic integrated circuit based on inverse-designed photonic components, IEEE Photonics Technology Letters 33, 1289 (2021).
- Siew et al. [2021] S. Y. Siew, B. Li, F. Gao, H. Y. Zheng, W. Zhang, P. Guo, S. W. Xie, A. Song, B. Dong, L. W. Luo, C. Li, X. Luo, and G.-Q. Lo, Review of silicon photonics technology and platform development, Journal of Lightwave Technology 39, 4374–4389 (2021).
- Logan et al. [2018] A. D. Logan, M. Gould, E. R. Schmidgall, K. Hestroffer, Z. Lin, W. Jin, A. Majumdar, F. Hatami, A. W. Rodriguez, and K.-M. C. Fu, 400W second harmonic conversion efficiency in 14 m-diameter gallium phosphide-on-oxide resonators, Optics Express 26, 33687 (2018).
- Guo et al. [2016] X. Guo, C.-l. Zou, C. Schuck, H. Jung, R. Cheng, and H. X. Tang, Parametric down-conversion photon-pair source on a nanophotonic chip, Light: Science Applications 6, e16249–e16249 (2016).
- Wang et al. [2023] X. Wang, X. Jiao, B. Wang, Y. Liu, X.-P. Xie, M.-Y. Zheng, Q. Zhang, and J.-W. Pan, Quantum frequency conversion and single-photon detection with lithium niobate nanophotonic chips, npj Quantum Information 9, 10.1038/s41534-023-00704-w (2023).
- Zhu et al. [2016] D. Zhu, H. Choi, T.-J. Lu, Q. Zhao, A. Dane, F. Najafi, D. R. Englund, and K. K. Berggren, Superconducting nanowire single-photon detector on aluminum nitride, in 2016 Conference on Lasers and Electro-Optics (CLEO) (2016) pp. 1–2.
- Esmaeil Zadeh et al. [2021] I. Esmaeil Zadeh, J. Chang, J. W. N. Los, S. Gyger, A. W. Elshaari, S. Steinhauer, S. N. Dorenbos, and V. Zwiller, Superconducting nanowire single-photon detectors: A perspective on evolution, state-of-the-art, future developments, and applications, Applied Physics Letters 118, 10.1063/5.0045990 (2021).
- Mouradian et al. [2015] S. L. Mouradian, T. Schröder, C. B. Poitras, L. Li, J. Goldstein, E. H. Chen, M. Walsh, J. Cardenas, M. L. Markham, D. J. Twitchen, M. Lipson, and D. Englund, Scalable Integration of Long-Lived Quantum Memories into a Photonic Circuit, Phys. Rev. X 5, 031009 (2015).
- Poberaj et al. [2012] G. Poberaj, H. Hu, W. Sohler, and P. Günter, Lithium niobate on insulator (LNOI) for micro-photonic devices, Laser Photonics Reviews 6, 488 (2012).
- Qi and Li [2020] Y. Qi and Y. Li, Integrated lithium niobate photonics, Nanophotonics 9, 1287–1320 (2020).
- Saravi et al. [2021] S. Saravi, T. Pertsch, and F. Setzpfandt, Lithium niobate on insulator: An emerging platform for integrated quantum photonics, Advanced Optical Materials 9, 10.1002/adom.202100789 (2021).
- Zhu et al. [2021] D. Zhu, L. Shao, M. Yu, R. Cheng, B. Desiatov, C. J. Xin, Y. Hu, J. Holzgrafe, S. Ghosh, A. Shams-Ansari, E. Puma, N. Sinclair, C. Reimer, M. Zhang, and M. Lončar, Integrated photonics on thin-film lithium niobate, Adv. Opt. Photonics 13, 242 (2021).
- Han et al. [2023] X. Han, M. Yuan, H. Xiao, G. Ren, T. G. Nguyen, A. Boes, Y. Su, A. Mitchell, and Y. Tian, Integrated photonics on the dielectrically loaded lithium niobate on insulator platform, Journal of the Optical Society of America B 40, D26 (2023).
- Zivasatienraj et al. [2021] B. Zivasatienraj, M. B. Tellekamp, and W. A. Doolittle, Epitaxy of LiNbO: Historical challenges and recent success, Crystals 11, 397 (2021).
- Buzaverov et al. [2023] K. A. Buzaverov, A. S. Baburin, E. V. Sergeev, S. S. Avdeev, E. S. Lotkov, M. Andronik, V. E. Stukalova, D. A. Baklykov, I. V. Dyakonov, N. N. Skryabin, M. Y. Saygin, S. P. Kulik, I. A. Ryzhikov, and I. A. Rodionov, Low-loss silicon nitride photonic ICs for near-infrared wavelength bandwidth, Optics Express 31, 16227 (2023).
- Hermans et al. [2019] A. Hermans, M. V. Daele, J. Dendooven, S. Clemmen, C. Detavernier, and R. Baets, Integrated silicon nitride electro-optic modulators with atomic layer deposited overlays, Opt. Lett. 44, 1112 (2019).
- Nejadriahi et al. [2020] H. Nejadriahi, A. Friedman, R. Sharma, S. Pappert, Y. Fainman, and P. Yu, Thermo-optic properties of silicon-rich silicon nitride for on-chip applications, Opt. Express 28, 24951 (2020).
- Li et al. [2021] N. Li, C. P. Ho, S. Zhu, Y. H. Fu, Y. Zhu, and L. Y. T. Lee, Aluminium nitride integrated photonics: a review, Nanophotonics 10, 2347–2387 (2021).
- Sun et al. [2019] Y. Sun, W. Shin, D. A. Laleyan, P. Wang, A. Pandey, X. Liu, Y. Wu, M. Soltani, and Z. Mi, Ultrahigh Q microring resonators using a single-crystal aluminum-nitride-on-sapphire platform, Optics Letters 44, 5679 (2019).
- Wu et al. [2020] X. Wu, J. Feng, X. Liu, , and H. Zeng, Effects of rapid thermal annealing on aluminum nitride waveguides, Optical Materials Express 10, 10.1364/OME.410129 (2020).
- Liu et al. [2018] X. Liu, A. W. Bruch, Z. Gong, J. Lu, J. B. Surya, L. Zhang, J. Wang, J. Yan, and H. X. Tang, Ultra-high-Q UV microring resonators based on a single-crystalline AlN platform, 5, 1279 (2018).
- Liu et al. [2017a] X. Liu, C. Sun, B. Xiong, L. Wang, J. Wang, Y. Han, Z. Hao, H. Li, Y. Luo, J. Yan, T. Wei, Y. Zhang, and J. Wang, Aluminum nitride-on-sapphire platform for integrated high-Q microresonators, Opt. Express 25, 587 (2017a).
- Hagedorn et al. [2019] S. Hagedorn, S. Walde, A. Mogilatenko, M. Weyers, L. Cancellara, M. Albrecht, and D. Jaeger, Stabilization of sputtered AlN/sapphire templates during high temperature annealing, Journal of Crystal Growth 512, 142 (2019).
- Tokumoto et al. [2009] Y. Tokumoto, N. Shibata, T. Mizoguhci, T. Yamamoto, and Y. Ikuhara, Atomic structure of threading dislocations in AlN thin films, Physica B: Condensed Matter 404, 4886 (2009).
- Dogheche et al. [2000] E. Dogheche, B. Belgacem, D. Remiens, P. Ruterana, and F. Omnes, Prism coupling as a non-destructive tool for optical characterization of (Al, Ga) nitride compounds, MRS Internet Journal of Nitride Semiconductor Research 5, 747 (2000).
- Schwarz et al. [2004] U. T. Schwarz, E. Sturm, W. Wegscheider, V. Kümmler, A. Lell, and V. Härle, Excitonic signature in gain and carrier induced change of refractive index spectra of (In, Al)GaN quantum well lasers, Applied Physics Letters 85, 1475 (2004).
- Bulutay et al. [2010] C. Bulutay, C. M. Turgut, and N. A. Zakhleniuk, Carrier-induced refractive index change and optical absorption in wurtzite InN and GaN: Full-band approach, Physical Review B 81, 10.1103/physrevb.81.155206 (2010).
- Udabe et al. [2023] A. Udabe, I. Baraia-Etxaburu, and D. G. Diez, Gallium nitride power devices: A state of the art review, IEEE Access 11, 48628 (2023).
- A. Tabbakh et al. [2023] T. A. Tabbakh, D. Anandan, M. J. Sheldon, P. Tyagi, and A. Alfaifi, Recent advancements in GaN LED technology, in Light-Emitting Diodes - New Perspectives (IntechOpen, 2023).
- Liang et al. [2021] F. Liang, D. Zhao, Z. Liu, P. Chen, J. Yang, L. Duan, Y. Shi, and H. Wang, GaN-based blue laser diode with 6.0 W of output power under continuous-wave operation at room temperature, Journal of Semiconductors 42, 112801 (2021).
- Hu et al. [2020] L. Hu, X. Ren, J. Liu, A. Tian, L. Jiang, S. Huang, W. Zhou, L. Zhang, and H. Yang, High-power hybrid GaN-based green laser diodes with ITO cladding layer, Photonics Research 8, 279 (2020).
- [37] F. D. Leonardis, R. A. Soref, M. Soltani, and V. M. N. Passaro, Broadband biphoton generation and statistics of quantum light in the UV-visible range in an AlGaN microring resonator 10.1038/s41598-017-11617-y.
- Leonardis et al. [2018] F. D. Leonardis, R. A. Soref, M. Soltani, and V. M. Passaro, Stimulated Brillouin scattering in an AlGaN photonics platform operating in the visible spectral range, Scientific Reports 8, 10.1038/S41598-018-33183-7 (2018).
- Li et al. [2004] N. Li, I. Waki, C. Kumtornkittikul, J. H. Liang, M. Sugiyama, Y. Shimogaki, and Y. Nakano, Fabrication of AlGaN-based waveguides by inductively coupled plasma etching, Japanese Journal of Applied Physics, Part 2: Letters 43, 10.1143/JJAP.43.L1340 (2004).
- Bruch et al. [2017] A. W. Bruch, K. Xiong, H. Jung, X. Guo, C. Zhang, J. Han, and H. X. Tang, Electrochemically sliced low loss AlGaN optical microresonators, Applied Physics Letters 110, 10.1063/1.4973521 (2017).
- Hagedorn et al. [2020] S. Hagedorn, S. Walde, A. Knauer, N. Susilo, D. Pacak, L. Cancellara, C. Netzel, A. Mogilatenko, C. Hartmann, T. Wernicke, M. Kneissl, and M. Weyers, Status and prospects of AlN templates on sapphire for ultraviolet light‐emitting diodes, physica status solidi (a) 217, 10.1002/pssa.201901022 (2020).
- Miyake et al. [2016] H. Miyake, C.-H. Lin, K. Tokoro, and K. Hiramatsu, Preparation of high-quality AlN on sapphire by high-temperature face-to-face annealing, Journal of Crystal Growth 456, 155 (2016).
- Pantha et al. [2007] B. N. Pantha, R. Dahal, M. L. Nakarmi, N. Nepal, J. Li, J. Y. Lin, H. X. Jiang, Q. S. Paduano, and D. Weyburne, Correlation between optoelectronic and structural properties and epilayer thickness of AlN, Applied Physics Letters 90, 10.1063/1.2747662 (2007).
- Weber et al. [1973] H. P. Weber, F. A. Dunn, and W. N. Leibolt, Loss measurements in thin-film optical waveguides, Appl. Opt. 12, 755 (1973).
- Dogheche et al. [1999] E. Dogheche, D. Rémiens, A. Boudrioua, and J. C. Loulergue, Growth and optical characterization of aluminum nitride thin films deposited on silicon by radio-frequency sputtering, Applied Physics Letters 74, 1209 (1999).
- Cannon et al. [2023] J. K. Cannon, S. G. Bishop, J. P. Hadden, H. B. Yağcı, and A. J. Bennett, Polarization study of single color centers in aluminum nitride, Applied Physics Letters 122, 10.1063/5.0145542 (2023).
- Varley et al. [2016] J. B. Varley, A. Janotti, and C. G. V. de Walle, Defects in AlN as candidates for solid-state qubits, Physical Review B 93, 10.1103/physrevb.93.161201 (2016).
- Xue et al. [2022] Y. Xue, J. Yuan, Q. Li, F. Chen, X. Yuan, Z. Ju, S. Zhang, B. Wu, Y. Hou, M. Li, J. Zhang, and E. Wu, Deciphering the photophysical properties of near-infrared quantum emitters in AlGaN films by transition dynamics, Nanoscale 14, 18115 (2022).
- Watanabe et al. [2008] N. Watanabe, T. Kimoto, and J. Suda, The temperature dependence of the refractive indices of GaN and AlN from room temperature up to 515 °C, Journal of Applied Physics 104, 10.1063/1.3021148 (2008).
- Zhou et al. [2022] H. Zhou, J. Dong, J. Cheng, W. Dong, C. Huang, Y. Shen, Q. Zhang, M. Gu, C. Qian, H. Chen, Z. Ruan, and X. Zhang, Photonic matrix multiplication lights up photonic accelerator and beyond, Light: Science Applications 11, 10.1038/s41377-022-00717-8 (2022).
- Clements et al. [2016] W. R. Clements, P. C. Humphreys, B. J. Metcalf, W. S. Kolthammer, and I. A. Walmsley, Optimal design for universal multiport interferometers, Optica 3, 1460 (2016).
- Liu et al. [2017b] Y. Liu, Q. X. Li, L. Y. Wan, B. Kucukgok, E. Ghafari, I. T. Ferguson, X. Zhang, S. Wang, Z. C. Feng, and N. Lu, Composition and temperature dependent optical properties of AlGaN alloy by spectroscopic ellipsometry, Applied Surface Science 421, 389 (2017b).
- Pernice et al. [2012] W. H. Pernice, C. Xiong, C. Schuck, and H. X. Tang, Second harmonic generation in phase matched aluminum nitride waveguides and micro-ring resonators, Applied Physics Letters 100, 10.1063/1.4722941 (2012).

### v2

@@ II.4 Tunable Microring Resonators
We fabricate microring resonators with a radius of 50 µm and a top width of 580 nm, coupled to racetrack waveguides with the same width and cleaved at both ends, as depicted in Fig. 4(a). The sample is mounted on a 3D piezoelectric nanopositioner (Physik Instrumente, NanoCube) and characterized with a custom-made inverted microscope, where the same objective (Zeiss, LD EC Epiplan-Neofluar, 20X, NA = 0.22) is used to launch the laser light into the input waveguide facet and collect it from the output one. The alignment of the incoupling light is monitored with an EMCCD camera (Andor, iXon Ultra 897) and optimized by maximizing the intensity of the outcoupled signal by adjusting both sample position and laser incoupling angle. Propagation losses are then inferred by measuring the intrinsic quality factor of the microring resonators. To reach critical coupling, which allows for the estimation of resonator losses, we fabricate several microring resonators coupled to racetrack waveguides with different coupling gaps, ranging from 100 to 325 nm, as displayed in Fig. 4(c). We measure the transmission of a tunable 785 nm laser (Toptica, DFB PRO centered at 784.6 nm, MHF over 2 nm) and use a linear film polarizer (Thorlabs, LPVIS) to excite and collect TE or TM modes selectively. Fig. 4(b) illustrates the typical transmission as a function of the laser’s frequency, with a baseline primarily attributed to the interference in measurement optics. This modulation can be described with a quadratic baseline from which the resonant peaks can be discriminated and fitted with a Lorentzian function, as indicated in red in Fig. 4(b). We extrapolate the resonance widths at the critical coupling, i.e., zero transmittance. The key parameters, such as the critical coupling FWHM, free spectral ranges (FSRs), and intrinsic Q-factors for TE and TM modes, are summarized in Table 1.
@@
| at 632 nm | 2.13 | 2.19 |
| at 632 nm | 2.05 | 2.10 |
We mount the chips on a custom-made temperature-controlled stage to test the temperature tunability of the TE and TM modes resonance frequencies in the range of 20-30 ∘C. Fig. 4(c) depicts the shift of a resonant peak of the TE mode (coupling gap=100nm) for different temperatures. In the investigated temperature range, the central frequency of the resonance changes linearly with temperature. The corresponding temperature coefficients are 5.7 and 6.2 GHz/K for TE and TM modes, respectively. These coefficients correspond to a thermo-optic coefficient of 2.9 10-5 K-1 and 3.6 10-5 K-1for TE and TM polarizations in the AlGaN waveguide, which is comparable to the previously reported values for AlN (2.94 10-5 K-1) and GaN (7.01 10-5 K-1) [56] at 785 nm and 300K.
II.5 Directional Couplers
@@
Directional couplers are essential components in photonic integrated circuits. They enable the transfer of electromagnetic energy between two or more waveguides that are in proximity. They can be combined to form optical switches or Mach-Zehnder interferometers, which are the fundamental constituents of most on-chip quantum technology applications [57, 58]. In our study, we develop AlGaN directional couplers linked to racetrack waveguides, as depicted in Fig. 5(a). We vary coupling gaps and lengths, as illustrated in Fig. 5(b). By injecting a 632 nm laser into one of the input ports, we measure the output intensity ratio: . Our observations reveal a direct linear correlation between the coupling length and , with linear fitting of the data providing the coupling constants. For coupling gaps of 230 nm and 330 nm, the coupling constants were determined to be 0.15 m-1 and 0.066 m-1, respectively. The observed uncertainties are primarily attributed to reflections at the input and output facets, which can be minimized by applying an anti-reflection coating. The well-functioning directional couplers show that our AlGaN platform is well-suited for building the required components of complex integrated photonic circuits. Further improvement is to be expected by optimizing the etching processes towards vertical sidewalls.
II.6 Phase Matching for Nonlinear Frequency Conversion and Pair Generation
In nonlinear photonics applications, achieving a phase matching of the modes involving the nonlinear conversion process is critical for efficiency. For integrated photonic waveguides, a common strategy is to match the effective indices of two distinct modes, typically a fundamental mode and a higher order mode [43, 59]. These approaches, however, suffer from a weak overlap between the involved modes. The mode overlap integral for second harmonic generation and SPDC for degenerate photon pairs is defined as [60]
where is the second-order nonlinear susceptibility tensor, and are the normalized mode electric field amplitudes of the modes at and , respectively. The integral is calculated over the mode volume . In the case of phase-matched TM modes for AlGaN waveguides, the dominating term in the overlap integral is
where is the relevant component of for the interaction along the primary axis of the crystal in TM modes, and and are the y-components of the electric field amplitudes. AlxGa1-xN alloys exhibit a sign change of the second-order nonlinear coefficient tensor component () at an alloy content of approximately = 0.65 [41]. We use this property to enhance the mode overlap. We design a heterostructure to achieve phase matching between the fundamental TM mode at 1550 nm and a higher-order mode at 775 nm with identical polarization, as demonstrated in Fig. 6(a-c). This strategy facilitates efficient phase matching, enabled by the sign change of , which significantly increases the overlap integral between the pump and SPDC modes. Our concept also allows for fine-tuning phase matching by adjusting the pump wavelength. The resultant effective index difference, , between the pump and down-conversion modes, which converge to zero at approximately 775 nm, is depicted in Fig. 6(d). For this simulation, we used a finite element solver (Comsol Multiphysics) to calculate the modes’ effective indices and used a genetic algorithm (supplied by Matlab) to minimize the objective function, . From bottom to top, optimized AlGaN layer thicknesses are 110/70/290/50 nm on 400 nm AlN on sapphire, while the alloy compositions are 0.55/0.42/0.65/0.66. The waveguide’s top width is 1.98 m, and the AlN layer is over-etched to a depth of 60 nm. Additionally, we define a 30 nm conformal AlN cladding. This particular layer design will likely exhibit some degree of strain relaxation, especially in the layers with low Al content. Changes in strain and piezoelectric polarization may impact the alignment of this calculation with experimental results. However, the overall viability of this method remains unaffected. The AlGaN/AlN heterostructure concept enables the engineering of compact SPDC sources and phase-matched ring resonators across a wide spectral range when coupled to a single-mode waveguide in a ring waveguide.
III Discussion
@@
We present a new material platform based on AlGaN heterostructures, which can be used to implement photonic devices with the core functionalities required to develop advanced photonic integrated chips. By epitaxially growing AlGaN on HTA-AlN templates on sapphire substrates, we alleviate the challenges related to crystal defects and dislocations at the sapphire-AlN interface and achieved an RMS surface roughness as low as 0.5 nm. The optical losses we quantify, 2.5 dB/cm for TE0 mode and 7.1 dB/cm for TM0 mode at 785 nm, outperform earlier measurements on aluminum nitride on sapphire waveguides in the visible spectrum. For instance, Lu et al. found losses of 5.3 dB/cm at 633 nm for the TE0 mode [61]. It is worth noting that we did not use a cladding material, which typically helps to reduce scattering loss further. By optimizing the layer sequence and growth recipes, as well as device fabrication, further reduction of losses will be possible. Moreover, by measuring the thermo-optic coefficient and anisotropic refractive indices of the waveguides, we have gained insight into parameters relevant to applications such as nonlinear photon-pair generation and on-chip nonlinear conversion. The birefringent refractive index of our devices opens avenues for exploration, especially in its utilization for phase matching of TE and TM modes. Furthermore, the surface roughness of the epitaxial material is sufficiently low to integrate superconducting single-photon detectors [62]. Our work introduces AlGaN heterostructures as a novel platform for integrated photonics, offering a scalable and high-quality alternative to well-established materials.
@@ Acknowledgements.
We thank Torsten Petzke and Cornelia Neumann for assistance in AlGaN growth, Karina Ickert and Nico Sluka for their help in spin-coating and lithography, Joost Wartena, Adrian Runge, Mohammad Mohammadi, and Dominik Sudau for their assistance in thin-film depositions; Kevin Kunkel, Natalia Sabelfeld, Kai Gehrke, and Andreas Renkewitz for performing plasma etching processes; Alexander Külberg and Felix Eiche for laser cutting and marking; Uwe Spengler and Christine Münnich for cleaving, Olaf Krüger, Ralph-Stephan Unger and Ina Ostermay for fruitful discussions. This work received support from the European Research Council (ERC), Starting Grant project QUREP, No. 851810, the German Federal Ministry of Education and Research (BMBF, project DiNOQuant, No. 13N14921; project QPIS, No. 16KISQ032K; project QPIC-1, No. 13N15858) and the Einstein Foundation Berlin (Einstein Research Unit: Perspectives of a quantum digital transformation: Near-term quantum computational devices and quantum processors).contributions
S.G. led microfabrication, characterization, simulations, and manuscript writing; S.P. specialized in ring resonator characterization; T.P. conducted e-beam lithography and supervised the fabrication processes; T.K. and S.H. were responsible for AlGaN material growth, M.W. and T.S. led the project. All authors reviewed the manuscript.